OpenStudy (anonymous):

7+52 to the power 2 *4

jhonyy9 (jhonyy9):

rewrite it please

OpenStudy (anonymous):

\[(7+52)^2\times4\]

OpenStudy (anonymous):

7 + 52power2 • 4

OpenStudy (anonymous):

is it so

OpenStudy (anonymous):

is giving me 472

OpenStudy (apoorvk):

\[(7 +52)^2 * 4\] \[=(59)^2 * 4\] \[=(60-1)^2 * 4\] \[=(3600 + 1 - 2*36 *1) * 4\] \[=(3601 - 72)*4\] \[=(3529)*4\] \[=14116\]

OpenStudy (anonymous):

no that cant not be the answer beacause in the worksheet i only have 214 35 128 107

OpenStudy (apoorvk):

If I misunderstood it, then it must be this \[7 + (50+2)^2*4\] \[=7 + (2500 + 4 + 100)*4\] \[=7 + 4(2604)\]

OpenStudy (apoorvk):

Then there are just two possibilities left 1. You posted it wrong. 2. Printing/typing error in your question or options

OpenStudy (anonymous):

this is the question on the worksheet evaluate 7+52 to the power 2 *4

OpenStudy (anonymous):

and the answer are 214 35 128 107

OpenStudy (anonymous):

\[7+52^{2}*4\]